# A Typology of Synthetic Datasets for Dialogue Processing in Clinical Contexts

### v2

@@ A Typology of Synthetic Datasets for
bedricks@ohsu.edu, as.dogruoz@ugent.be, sergiu.nisioi@unibuc.ro
@@ 1.  Introduction
A particularly fast-growing and relevant domain of clinical language is that of the “clinical dialogue (i.e., a dialogue between a patient and a health provider which can also be referred to as a clinical or medical conversation). NLP tools to process such language are seeing rapid adoption in practice, most notably in the form of tools for automated generation of clinical documentation (Tierney:2024; Finley:2018) and for more traditional dialogue applications (e.g., clinical chatbots). Such dialogues may occur via spoken language, or may be written (taking place via an online discussion forum, a chat-style interface, etc.).
Because of the intense commercial focus in this space, there exists a high level of need for representative dialogue datasets for system development and evaluation. However, due to extremely sensitive nature of clinical conversations, genuine data is very difficult to obtain, even by the standards of clinical NLP. As such, there are a wide range of synthetic datasets purporting to represent clinical conversations of various types.
While synthetic datasets have been shown to be sufficient in some situations, little theory exists to inform the scientific audience about how these datasets may be best used and generalized to new NLP applications. Towards that end, our goals in this paper are 1) to provide an overview of how synthetic data sets are created, evaluated and used for dialogue-related tasks in the medical domain via a literature review, and 2) to classify these data sets according to a novel typology.
One of the challenges toward this goal is a lack of consensus on what is meant by the word “synthetic” in terms of datasets on dialogues. The canonical definition of synthetic data is that of the census:2025aa, which may be summarized as follows: given a dataset $D$, generate a new dataset $D^{\prime}$ in a way that a) individual instances in $D^{\prime}$ are not present in $D$, b) the relevant statistical properties of $D^{\prime}$ match those of $D$, and c) original instances from $D$ may not be recovered from $D^{\prime}$. [^1] This formulation echos that proposed in the earliest literature on synthetic microdata (DonaldB.Rubin:1993). [^2] Throughout this paper, we use “microdata” in the statistical sense of the term; i.e., referring to the individual data points contained in the dataset: sentences in a corpus, images in a collection, responses from a survey, etc. This definition fits well in the context of a dataset consisting of structured data. However, as noted out by aggarwalGeneralSurveyPrivacyPreserving2008, it is more challenging to apply this to textual data, for three primary reasons.
@@
Another challenging aspect of synthetic data is that artificiality is not a binary property. In that sense, a dataset may be “more” or “less” synthetic, depending on its construction and the point of view of evaluation. All datasets, being curated and abstracted representations of some sampled reality, are inherently synthetic to some degree; this process necessarily involves intervention around questions of inclusion/exclusion, representation, and structure. As bowkerMemoryPracticesSciences2005 put it, “raw data is both an oxymoron and a bad idea; to the contrary, data should be cooked with care.” However, while all datasets may be synthetic, some datasets are more synthetic than others. Literature on synthetic data distinguishes between datasets that are “fully synthetic,” being those that contain no original, genuine microdata, and “partially synthetic,” in which only certain variables are perturbed reiterInferencePartiallySynthetic2003. In terms of structured medical data, Gonzales:2023 and Murtaza:2023 provide comprehensive discussions on this and other relevant taxonomies. How can this framework, originally designed for numerical data, be applied to linguistic data, which focuses on meaning and grammar rather than statistics?
@@ 2.  Means of Production
Synthetic data production methods are frequently categorized into “process-driven” [^3] Sometimes referred to as “theory-” or “knowledge-driven.” and “data-driven” methods. Under the former, whatever generative process is carrying out the synthesis is informed by human knowledge and/or by information that is exogenous to the dataset itself. The latter describes methods that use properties of the dataset itself to derive the generative model (Goncalves:2020a; Murtaza:2023).
This framing emphasizes description of the mechanical process by which the synthesis is done. While important, it does not describe what is being synthesized, nor does it address the degree to which a given dataset should be considered synthetic. This omission is problematic when considering synthetic linguistic data, as there are numerous aspects of the data that might (or might not) be synthesized: topic, genre/register, grammatical or discursive structure, specific aspects of semantic content, individual lexical choices, etc. A given dataset might involve synthesis along one of these domains, while leaving other domains un-perturbed. To illustrate this issue, consider the question of how synthetic clinical datasets are created.
@@
In simulated scenarios involving humans, the pragmatics and paralinguistic components of resulting dialogues may be less representative of what would be seen in a true naturalistic exchange. Similarly, we would expect a “screenwritten” scenario imagined by a single author to be less representative than a simulation involving multiple humans interacting in a more naturalistic way. However, depending on the intended use of the resulting data, this may not be an issue: for example, if the primary content of interest is the semantic or lexical contents of the exchange, as opposed to the pragmatic or discursive contents. [^4] See section 4.2 for additional discussion.
Finally, simulated datasets may be created mechanically de novo using a natural language generation (NLG) model (frei2023annotated). This may take the form of prompted generation (i.e., an LLM being prompted to write a note or simulate a dialogue given certain criteria) or in some cases by literally having two LLMs “talk” to one another in a simulated dialogue, in an effort to increase the variety of the resulting output (grayIncreasingRealismVariety2024).
@@
Similarly, we argue that an otherwise-genuine dataset that has been “anonymized” is, in some sense synthetic (the degree to which this is the case depends on the method of anonymization in question). At one end of the spectrum, imagine a scheme in which identifiers are replaced with a class token of some kind (e.g. PATIENT_NAME, DAY_OF_WEEK, etc.) in a purely deterministic and uniform manner. At the other end of that spectrum, consider a method in which identifiers are replaced with statistically-plausible alternative values (“pseudonymization”) [^5] See uzunerEvaluatingStateoftheartAutomatic2007 for a detailed description of such a method, as well as an overview on deidentification in non-dialogue clinical text in general.
In both cases, the process of applying such a method results in the creation of a new and distinct dataset. We argue, however, that the second case is “more synthetic” in that there are two levels of synthesis in play: first, the simple alteration of the original un-modified data, and second, in that some designed process informed the pseudonymization. This process must itself necessarily bring with it working assumptions, and may also depend on yet additional sources of data, further distancing the resulting synthetic dataset from whatever grounding in reality it may have had earlier in its development.
@@ 3.  What Do We Mean By “Synthetic”?
In response to the complexities surrounding the question of whether a given dataset is “synthetic” or not, we propose a conceptual framework whose aim is to provide structure and clarity when describing datasets that have a synthetic component, and to facilitate meaningful comparisons. For our purposes, we define a dataset as a collection of information-bearing objects that have been purposefully assembled or otherwise grouped, with the goal of serving a particular representational function, with the intention of being used in analysis (broadly defined) of some kind.
@@
How does our above definition fit with synthetic data? A dataset is, per definition, a human-generated artifact. At all stages of collection, curation, and processing, choices are made that separate the resulting dataset from its originating context. As such, we further define a “naturalistic” dataset as one that fulfills the terms of our above definition, along with the additional constraint that the microdata contained in the dataset are a) intended by the designers to be faithful representations of some genuine originating phenomenon, and b) are understood to be such by the users of the dataset.
@@
For example, consider the MedDialog corpus of patient-provider dialogues taken from an online medical consultation website (zeng-etal-2020-meddialog). The designers of this dataset understood and believed these dialogues to represent exchanges that actually took place between human patients and human providers who actually existed. While they may have applied criteria to which dialogues were included (perhaps focusing on a particular category of patient, or clinical specialty), the lexical, semantic, and pragmatic properties of the language were left un-modified in the hope of comprising a maximally representative sample. In turn, downstream users of this dataset would understand its contents to be representative of exchanges that actually took place (i.e., to be “real messages”).
@@
In comparison, consider the aci subset of the Aci-bench dataset (yim2023aci), which consists of transcribed exchanges between an actual doctor and a volunteer “patient,” acting out a scenario based on symptom-driven prompts. The exchanges in such a dataset are “real” in the sense that they did, in fact, take place, and occurred between actual humans. However, they are not “real” in that they arose from an artificial scenario, as opposed from the actual real-world domain that they are attempting to represent.
@@
The creators of Aci-bench, of course, attempted to design their scenarios to be as “realistic” as possible, and gave careful consideration to their design. Their goal was to create a dataset whose synthetic features did not affect those aspects of its form or content to an extent that compromised the dataset’s fundamental validity [^6] Here we use “validity” in the specific senses of construct and measurement validity; see brewer_crano_2014 for an overview of these terms, and jacobs_wallach_2021 for a discussion of how these concepts apply in machine learning research. for the analyses of the sort that the dataset was intended by its authors and users to support. Decisions about whether, when, and how to use such a dataset typically hinge on the extent to which this assumption holds for one’s particular purpose.
@@
Finally, consider the DoPaCo dataset  (chen2023investigating). Here, the authors began with transcribed and de-identified genuine doctor-patient conversations, to which they had extremely restricted and time-limited access, and which could not be re-distributed. They used this corpus to fine-tune a pre-trained dialogue-generation Transformer model, and then used this model to generate a large corpus of synthetic dialogues, which they used for the rest of their work. One may argue that such a dataset is in some sense meaningfully or usefully representative of a “real” dataset of doctor-patient conversations, according to any number of rubrics. [^7] Given that it was produced via a sequence-to-sequence model, one may accurately and literally describe it as consisting of a set of samples drawn from a statistical model of the originating “real” dataset; as such, there exist statistical tools that we may use to quantify the degree to which such samples are or are not representative (in a purely statistical sense) of the originating data, such as Perplexity and KL-Divergence. However, we note that there is a long-standing and robust literature (Galliers:1993; iyer_1997; ito_1999) on the limitations of such purely statistical measures in terms of their external and ecological validity (brewer_crano_2014; Egami:2023). But, it is also equally true that such a dataset, however statistically similar it may be to its original training data, is not comprised of “real” dialogue data in the same sense as Aci-bench or MedDialog. Rather than representing true discursive choices made by humans interacting with one another, the dialogues in the DoPaCo dataset reflect that model’s statistical approximation of such choices, as represented its training data.
@@
This question of “real-ness” is an entirely separate question from whether such a dataset is useful or sufficient for a given purpose. We are again using “real” as a descriptive rather than a normative term, and of course even highly synthetic datasets have many uses. As such, we are emphatically not arguing that datasets that are more naturalistic are inherently superior in some way to datasets that are more synthetic, though there may indeed exist scenarios where this is the case. However, that question is orthogonal to our purpose in the present work. Finally, we note that the synthetic datasets we describe in this paper are accompanied by more or less detailed evaluations by their creators, seeking to investigate the degree to which they are or are not sufficient for specific applications.
@@ 3.1.  A Typology for Human/Machine Microdata Interventions
However, if the point of the dataset was instead to enable research into the quality or utility of human-written dialogues. In other words, if the dataset was made from simulated dialogues, we would instead consider the human-written dialogues to be non-synthetic primary microdata, and categorize the human component of this dataset under “Type 1.”
@@ 3.2.  In Search of Synthetic Dialogues
Included publications could be specifically focused on the production and evaluation of their datasets, or could produce datasets as secondary research artifacts. In the latter case, we required that there be an intention on the part of the authors that their datasets be re-used by others (i.e., their data were not simply made re-usable for replicability). We used keyword-based and index-term search strategies, supplemented by large language models (LLMs) to assist in filtering and classifying retrieved papers.
@@
Given our focus on dialogues, we excluded the following data sets from our analyses in this paper:
@@
3. Datasets that use genuine dialogues, but whose artifacts of interest and analysis are synthetically-produced secondary products (such as summaries or clinical notes ) generated automatically from those original genuine dialogues, .
@@ 4.1.  Synthetic Data and Varieties of Task
As discovered in our review, synthetic data sets in clinical domain are used for a variety of NLP related tasks such as: Information Extraction (zhang-etal-2020-mie), summarization of the clinical dialogues (joshi-etal-2020-dr; chintagunta-etal-2021-medically), generating notes from conversations yim-etal-2020-alignment, dialogue with custom personality traits for training purposes han-etal-2024-psydial-personality, training data for ASR systems fareez2022dataset, and named entity recognition frei2023annotated.
@@
Type of Medical Domains: Most data sets we were able to retrieve are not targeted for a specific type of disease in the clinical domain, and instead attempted to more represent written clinical notes or patient-provider dialogues from generic clinical settings (e.g. an outpatient visit). However, we did encounter a handful of synthetic data sets focusing on specific clinical specialties, including cardiology (zhang-etal-2020-mie), gastroenterology (liu2022meddg), psychology (liu-etal-2021-towards; zheng-etal-2023-augesc; han-etal-2024-psydial-personality), and obstetrics & gynecology (Wang:2023d).

### v3

@@ A Typology of Synthetic Datasets for
bedricks@ohsu.edu, as.dogruoz@ugent.be, sergiu.nisioi@unibuc.rothanks: All authors are corresponding authors ♣♡♠.
@@ 1.  Introduction
A particularly fast-growing and relevant domain of clinical language is that of the “clinical dialogue” (i.e., a dialogue between a patient and a health provider, sometimes also referred to as a clinical or medical “conversation”). NLP tools to process such language are being rapidly adopted into practice, most notably in the form of tools for automated generation of clinical documentation (Tierney:2024; Finley:2018) as well as for more traditional dialogue applications (e.g., clinical chatbots). Such dialogues may occur via spoken language, or may be written (taking place via an online discussion forum, a chat-style interface, etc.).
Because of the intense commercial focus in this space, there exists a high level of need for representative dialogue datasets for system development and evaluation (dogruoz-etal-2023-representativeness; adelani-etal-2025-generative). However, due to the extremely sensitive nature of clinical conversations, genuine data is very difficult to obtain, even by the standards of clinical NLP. As such, there are a wide range of synthetic datasets purporting to represent clinical conversations of various types.
While synthetic datasets have been shown to be sufficient in some situations, little theory exists to inform scientific audiences about how these datasets may be best used and generalized to new NLP applications. Towards that end, our goals in this paper are: 1) to provide an overview of how such synthetic datasets are created, evaluated, and used for dialogue-related tasks in the medical domain via a literature review; and 2) to classify these datasets according to a novel typology.
One of the challenges toward this goal is a lack of consensus on what is meant by the word “synthetic” in terms of datasets on dialogues. The canonical definition of synthetic data is that codified by the U.S. Census Bureau in their Statistical Quality Standards (census:2025aa), which may be summarized as follows: given a dataset $D$, generate a new dataset $D^{\prime}$ in a way that a) individual instances in $D^{\prime}$ are not present in $D$, b) the relevant statistical properties of $D^{\prime}$ match those of $D$, and c) original instances from $D$ may not be recovered from $D^{\prime}$. [^1] This formulation echos that proposed in the earliest literature (DonaldB.Rubin:1993) on synthetic microdata, i.e., the individual data points contained in a dataset: sentences in a corpus, images in a collection, responses from a survey, etc. This definition fits well in the context of a dataset consisting of structured data. However, as noted by aggarwalGeneralSurveyPrivacyPreserving2008, it is more challenging to apply this to textual data, for three primary reasons.
@@
Another challenging aspect of synthetic data is that artificiality is not a binary property. All datasets, being curated and abstracted representations of some sampled reality, are inherently synthetic to some degree. This process necessarily involves intervention around questions of inclusion/exclusion, representation, and structure. As bowkerMemoryPracticesSciences2005 put it, “raw data is both an oxymoron and a bad idea; to the contrary, data should be cooked with care.” However, while all datasets may be synthetic, some datasets are more synthetic than others: a dataset may be “more” or “less” synthetic, depending on its construction and the point of view of evaluation. Literature on synthetic data distinguishes between datasets that are “fully synthetic,” being those that contain no original, genuine microdata, and “partially synthetic,” in which only certain variables are perturbed (reiterInferencePartiallySynthetic2003). In terms of structured medical data, Gonzales:2023 and Murtaza:2023 provide comprehensive discussions on this and other relevant taxonomies.
How can this framework, originally designed for numerical data, be applied to linguistic data, which focuses on meaning and grammar rather than statistics?
@@ 2.  Means of Production
Synthetic data production methods can be categorized into “process-driven” [^2] Sometimes referred to as “theory-” or “knowledge-driven.” and “data-driven” methods. Under the former, whatever generative process is carrying out the synthesis is informed by human knowledge and/or by information that is exogenous to the dataset itself. The latter describes methods that use properties of the dataset itself to derive the generative model (Goncalves:2020a; Murtaza:2023).
This framing emphasizes a description of the mechanical process by which the synthesis is done. While important, it does not describe what is being synthesized, nor does it address the degree to which a given dataset should be considered synthetic. This omission is problematic when considering synthetic linguistic data, as there are numerous aspects (e.g., topic, genre/register, grammatical or discursive structure, specific aspects of semantic content, individual lexical choices, etc.) of the data that might (or might not) be synthesized. A given dataset might involve synthesis along one of these domains, while leaving other domains un-perturbed. To illustrate this issue, consider the question of how synthetic clinical conversational datasets are created.
@@
In simulated scenarios involving humans, the pragmatic and paralinguistic components of the resulting dialogues may be less representative of what would be seen in a true naturalistic exchange. Similarly, we would expect a “screenwritten” scenario imagined by a single author to be less representative than a simulation involving multiple humans interacting in a more naturalistic way. However, depending on the intended use of the resulting data, this may not be an issue. For example, if the primary content of interest is the semantic or lexical contents of the exchange, as opposed to the pragmatic or discursive contents, such a dataset may be entirely adequate. If, however, the analytical question of interest depends on discourse features, a synthetic dataset of this sort would likely be insufficiently representative. [^3] See section 4.2 for additional discussion.
Finally, simulated datasets may be created de novo via a mechanical process, e.g. using a natural language generation (NLG) model (frei2023annotated). This may take the form of prompted generation (i.e., an LLM being prompted to write a note or simulate a dialogue given certain criteria) or in some cases by literally having two LLMs “talk” to one another in a simulated dialogue, in an effort to increase the variety of the resulting output (grayIncreasingRealismVariety2024).
@@
Similarly, we argue that an otherwise-genuine dataset that has been “anonymized” is, in some sense synthetic, with the degree to which this is the case depending on the method of anonymization in question. At one end of the spectrum, imagine a scheme in which identifiers are replaced with a class token of some kind (e.g. PATIENT_NAME, DAY_OF_WEEK, etc.) in a purely deterministic and uniform manner. At the other end of that spectrum, consider a method in which identifiers are replaced with statistically-plausible alternative values (“pseudonymization”) [^4] See uzunerEvaluatingStateoftheartAutomatic2007 for a detailed description of such a method, as well as an overview on deidentification in non-dialogue clinical text in general.
In both cases, the process of applying such a method results in the creation of a new and distinct dataset. We argue, however, that the second case is “more synthetic” in that there are two levels of synthesis in play: the simple alteration of the original un-modified data, in combination with some designed process informing the pseudonymization. This process must itself necessarily bring with it working assumptions, and may also depend on yet additional sources of data, further distancing the resulting synthetic dataset from whatever grounding in reality it may have had earlier in its development.
@@ 3.  What Do We Mean By “Synthetic”?
In response to the complexities surrounding the question of whether a given dataset is “synthetic” or not, we propose a conceptual framework whose aim is to provide structure and clarity when describing datasets that have a synthetic component, and to facilitate meaningful comparisons. For our purposes, we define a dataset as a collection of information-bearing objects that have been purposefully assembled or otherwise grouped, with the goal of serving a particular representational function, with the intention of being used in analysis (broadly defined) of some kind. See Appendix A for a detailed discussion of this definition and its constituent elements.
@@
How does our above definition fit with synthetic data? A dataset is, by definition, a human-generated artifact. At all stages of collection, curation, and processing, choices are made that separate the resulting dataset from its originating context. As such, we further define a “naturalistic” dataset as one that fulfills the terms of our above definition, along with the additional constraint that the microdata contained in the dataset are a) intended by the designers to be faithful representations of some genuine originating phenomenon, and b) are understood to be such by the users of the dataset.
@@
For example, consider the MedDialog corpus of patient-provider dialogues taken from an online medical consultation website (zeng-etal-2020-meddialog). The designers of this dataset understood and believed these dialogues to represent exchanges that actually took place between human patients and human providers who actually existed. While they may have applied criteria to which dialogues were included (perhaps focusing on a particular category of patient, or clinical specialty), the lexical, semantic, pragmatic properties of the language and sociolinguistic aspects of the clinical dialogue were left un-modified in the hope of comprising a maximally representative sample. In turn, downstream users of this dataset would understand its contents to be representative of exchanges that actually took place (i.e., to be “real messages”).
@@
In comparison, consider the ACI subset of the ACI-bench dataset (yim2023aci), which consists of transcribed exchanges between an actual doctor and a volunteer “patient,” acting out a scenario based on symptom-driven prompts. The exchanges in such a dataset are “real” in the sense that they did, in fact, take place, and occurred between actual humans. However, they are not “real” in that they arose from an artificial scenario, as opposed to the actual real-world domain that they are attempting to represent.
@@
The creators of ACI-bench, of course, attempted to design their scenarios to be as “realistic” as possible, and gave careful consideration to their design. Their goal was to create a dataset whose synthetic nature did not affect those aspects of its form or content to an extent that compromised the dataset’s fundamental validity [^5] Here we use “validity” in the specific senses of construct and measurement validity; see brewer_crano_2014 for an overview of these terms, and jacobs_wallach_2021 for a discussion of how these concepts apply in machine learning research. for the analyses of the sort that the dataset was intended by its authors and users to support. Decisions about whether, when, and how to use such a dataset typically hinge on the extent to which this assumption holds for one’s particular purpose.
@@
Finally, consider the DoPaCo dataset  (chen2023investigating). Here, the authors began with transcribed and de-identified genuine doctor-patient conversations, to which they had extremely restricted and time-limited access, and which could not be re-distributed. They used this corpus to fine-tune a pre-trained dialogue-generation Transformer model, and then used this model to generate a large corpus of synthetic dialogues, which they used for the rest of their work. One may argue that such a dataset is in some sense meaningfully or usefully representative of a “real” dataset of doctor-patient conversations, according to any number of rubrics. [^6] Given that it was produced via a sequence-to-sequence model, one may accurately and literally describe it as consisting of a set of samples drawn from a statistical model of the originating “real” dataset; as such, there exist statistical measures that we may use to quantify the degree to which such samples are or are not representative (in a purely statistical sense) of the originating data, such as perplexity and KL-Divergence. However, we also note that there is a long-standing and robust literature (Galliers:1993; iyer_1997; ito_1999) on the limitations of such purely statistical measures in terms of their external and ecological validity (brewer_crano_2014; Egami:2023). But, it is also equally true that such a dataset, however statistically similar it may be to its original training data, is not comprised of “real” dialogue data in the same sense as ACI-bench or MedDialog. Rather than representing true discursive choices made by humans interacting with one another, the dialogues in the DoPaCo dataset reflect that model’s statistical approximation of the discursive choices represented in its training data.
@@
This question of “real-ness” is an entirely separate question from whether such a dataset is useful or sufficient for a given purpose. We are again using “real” as a descriptive rather than a normative term, and of course even highly synthetic datasets have many uses. As such, we are emphatically not arguing that datasets that are more naturalistic are inherently and categorically superior in some way to datasets that are more synthetic, though there may indeed exist scenarios where this is the case; that question is orthogonal to our purpose in the present work. Finally, we note that the synthetic datasets we describe in this paper are uniformly accompanied by more or less detailed evaluations by their creators, seeking to investigate the degree to which they are or are not sufficient for specific applications.
@@ 3.1.  A Typology for Human/Machine Microdata Interventions
However, if the goal of the dataset had been to consist of simulated dialogues (as opposed to the actual goal, to wit, using simulated dialogues as representative stand-ins for genuine dialogues), we would instead consider the human-written dialogues to be non-synthetic primary microdata, and categorize the human component of this dataset under “Type 1.”
@@ 3.2.  In Search of Synthetic Dialogues
Included publications could be specifically focused on the production and evaluation of their datasets, or could produce datasets as secondary research artifacts. In the latter case, we required that there be an intention on the part of the authors that their datasets be re-used by others (i.e., their data were not simply made re-usable for purposes of replicability). We used keyword-based and index-term search strategies, supplemented by large language models (LLMs) to assist in filtering and classifying retrieved papers. See Appendix B for a detailed description of our approach and Appendix D for the verbatim LLM prompts. Table 2 summarizes the results of our structured search strategies across PubMed, the ACL Anthology, and dblp.
@@
Given our focus on dialogues, we excluded the following categories of dataset from our analyses in this paper:
@@
3. Datasets that use genuine dialogues, but whose artifacts of interest and analysis are synthetically-produced secondary products (such as summaries or clinical notes) generated automatically from those original genuine dialogues.
@@ 4.1.  Synthetic Data and Varieties of Task
As discovered in our review, synthetic datasets in the clinical domain are used for a variety of NLP related tasks such as: Information Extraction (zhang-etal-2020-mie), summarization of the clinical dialogues (joshi-etal-2020-dr; chintagunta-etal-2021-medically), generating notes from conversations yim-etal-2020-alignment, dialogue with custom personality traits for training purposes han-etal-2024-psydial-personality, training data for ASR systems fareez2022dataset, and named entity recognition frei2023annotated.
@@
Type of Medical Domains: Most datasets we were able to retrieve are not targeted for a specific type of disease in the clinical domain, and instead attempted to more broadly represent written clinical notes or patient-provider dialogues from generic clinical settings (e.g. an outpatient visit). However, we did encounter a handful of synthetic datasets focusing on specific clinical specialties, including cardiology (zhang-etal-2020-mie), gastroenterology (liu2022meddg), psychology (liu-etal-2021-towards; zheng-etal-2023-augesc; han-etal-2024-psydial-personality), and obstetrics & gynecology (Wang:2023d).